# ClarifyCoder: Clarification-Aware Fine-Tuning for Programmatic Problem Solving

### v1

@@ Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities in code generation tasks. However, a significant gap remains between their current performance and that of expert software engineers. A key differentiator is that human engineers actively seek clarification when faced with ambiguous requirements, while LLMs typically generate code regardless of uncertainties in the problem description. We present ClarifyCoder, a novel framework with synthetic data generation and instruction-tuning that enables LLMs to identify ambiguities and request clarification before proceeding with code generation. While recent work has focused on LLM-based agents for iterative code generation, we argue that the fundamental ability to recognize and query ambiguous requirements should be intrinsic to the models themselves. Our approach consists of two main components: (1) a data synthesis technique that augments existing programming datasets with scenarios requiring clarification to generate clarification-aware training data, and (2) a fine-tuning strategy that teaches models to prioritize seeking clarification over immediate code generation when faced with incomplete or ambiguous requirements. We further provide an empirical analysis of integrating ClarifyCoder with standard fine-tuning for a joint optimization of both clarify-awareness and coding ability. Experimental results demonstrate that ClarifyCoder significantly improves the communication capabilities of Code LLMs through meaningful clarification dialogues while maintaining code generation capabilities.
@@ 1. Introduction
Large language models (LLMs) (Vaswani et al., 2017; Svyatkovskiy et al., 2020; Wang et al., 2021; Feng et al., 2020) have revolutionized problem-solving activities, such as code generation from natural language requirements, as demonstrated by systems like Codex (Chen et al., 2021), AlphaCode (Li et al., 2022a), and CodeGen (Nijkamp et al., 2022). While these advances demonstrate remarkable technical proficiency, critical limitations persist in their practical application to real-world problem-solving scenarios. Current models frequently exhibit issues such as intent misinterpretation due to ambiguous user requirements (Mu et al., 2024), insufficient problem decomposition strategies (Sarkar et al., 2022; Lahiri et al., 2022), hallucination of non-existent constraints, unidirectional communication patterns (Wu and Fard, 2025), and miscalibrated confidence levels that persist despite uncertain inputs (Liu et al., 2023b, a). These limitations reveal a fundamental disconnect between current AI coding assistants and the collaborative dynamics of human pair programming (Hassan et al., 2024).
Figure 1. Illustration of data construction of ClarifyCoder.
In real-world problem-solving processes, people, including software engineers, rely on continuous dialog to resolve ambiguities, align mental models, and iteratively refine problem statements and their solutions (Whitehead, 2007; Pressman, 2005; Mistrík et al., 2010). This communication-first paradigm—manifested through requirements clarification, design rationale discussions, and solution validation—forms the bedrock of reliable software development (McChesney and Gallagher, 2004; Jazayeri, 2004). This reflects a fundamental difference between natural language and programming language – natural language inherently contains ambiguities, while programming languages are designed to be precise and unambiguous (Hoare, 1969). Current academic benchmarks, however, mainly feature well-specified tasks with minimal real-world ambiguity (Allamanis et al., 2021), creating a false equivalence between exam-style performance and practical applicability. In production environments, 72% of software defects originate from misunderstood requirements (Group et al., 2015), underscoring the critical need for clarification-driven development processes. Although there has been recent research to explore the application of LLMs in requirements engineering (Guo et al., 2023; Zhao et al., 2023; Wang et al., 2023), in this work, we focus specifically on code generation tasks to explore how well LLM can ask necessary and high-quality clarifying questions from coding problem statements. However, despite this clarification-driven reality in the real-world, recent empirical studies reveal that state-of-the-art Code LLMs generate code outputs in over 63% of ambiguous scenarios without seeking necessary clarifications (Wu and Fard, 2025).
Addressing the above gap presents several key challenges. First, while recent approaches have focused on using LLM-based agents for iterative code generation (Mu et al., 2024; Miao et al., 2025), we argue that the ability to identify ambiguity and ask clarifying questions should be an intrinsic capability of the models themselves, which needs to be achieved by aligning models. This fundamental capability would enhance not just code generation but a broad range of software engineering tasks and problem-solving tasks. Second, existing instruction-tuning methods inadvertently reinforce the tendency to generate code regardless of input quality, similar to hallucination issues in natural language processing (Zhang et al., 2024). This limitation likely stems from the nature of current instruction tuning datasets, which assume that instructions are always precise enough. However, this assumption does not hold in real-world scenarios, where requirements are often ambiguous or incomplete.
To overcome these challenges, there is a critical need to create datasets that can both benchmark and improve LLMs’ requirement clarification capabilities. Such datasets would enable models to learn when and how to ask good questions, a skill that is currently underdeveloped due to the lack of appropriate training data. Finally, while datasets like HumanEvalComm (Wu and Fard, 2025) provide 164 human-curated examples of clarification dialogues, the size is still too small for conducting fine-tuning, and manually creating such datasets at scale is prohibitively expensive. This underscores the need for approaches to generate larger-scale, high-quality datasets that capture the nuances of requirement clarification in software engineering contexts.
@@
In this research, we present ClarifyCoder, a novel framework that includes synthetic data generation and instruction-tuning techniques to enable LLMs to recognize ambiguity and request clarification when needed. The main focus of our study is to fine-tune Code LLMs to develop the capability of asking effective clarifying questions. With the goal to enhance LLMs’ clarification awareness in programmatic problem-solving, the framework includes 1) a new synthetic data generation approach for automatically creating ambiguous problem descriptions and corresponding clarifying questions, and 2) a technique combining automated data synthesis with targeted instruction tuning for clarification-aware (or clarify-aware) code generation. As a result, ClarifyCoder can be integrated in the standard fine-tuning that satisfies both clarify-awareness and code generation capability.
Our study is motivated by the observation that current instruction tuning methods in Code LLMs force models to generate code regardless of issues in the problem description. This tendency can lead to incorrect or inappropriate code generation when faced with ambiguous or incomplete problem statements. We evaluate both the utility and clarify-awareness of our clarify-aware fine-tuning based on the proposed synthetic training data generation. Our experiments demonstrate that ClarifyCoder significantly improves LLMs’ ability to recognize ambiguity and generate appropriate clarifying questions, including increases in communication rate and good question rate by absolute 40% and 30%, respectively, while maintaining the code generation ability.
To summarize, our key contributions of ClarifyCoder include:
  • A new synthetic data generation approach for automatically generating ambiguous problem descriptions and corresponding clarifying questions.
@@
  • A clarify-aware framework that combines synthetic data generation with targeted fine-tuning to enable clarify-aware code generation.
  • Comprehensive evaluation and empirical study across multiple LLMs demonstrating ClarifyCoder’s effectiveness in improving clarify-aware capabilities.
@@
  • A publicly available, open-source repository (upon acceptance) for experimental reproduction and future research on ClarifyCoder, at https://github.com/jie-jw-wu/clarify-coder.
@@ 2.1. Backgrounds
  • Maintain the model’s code generation capabilities for well-specified tasks
  • Enable the model to recognize ambiguity and request clarification when necessary
@@
Table 1. Comparison of responses of DeepSeek Coder and ClarifyCoder for ambiguous problem statement. The ClarifyCoder is fine-tuned based on the DeepSeek Coder that shows its ability to ask clarifying questions to disambiguate coding problem. This is an incomplete problem in HumanEvalComm, modified from HumanEval/43 in HumanEval benchmark.
@@
[…]
@@ 2.2. Problem Formalization
Let $H$ donates the solution space that comprises all syntactically valid implementations :
@@
Definition 2.1 (Ambiguous Problem).
A coding problem $S$ exhibits ambiguity if $H^{*}\subset H$, i.e.,
@@
”””Return list with elements incremented by a
number.”””
This specification permits multiple valid interpretations:
$H$ =$=\{h_{\text{1}},h_{\text{2}},h_{\text{3}},h_{\text{4}},...\}$
[…]
In this example, the ambiguity stems from underspecified number to increment by. See also another example in Figure 2. Given this ambiguous problem, the goal of this model is to ask clarifying questions to disambiguate the coding problem. When the clarifying question(s) are answered, with this additional information, the model attempts to generate the correct code $h\in H^{*}$.
@@ 3.1. Baseline Instruction Tuning
We first establish a baseline by employing standard instruction tuning on the original programming dataset, denoted as $D^{og}$ This dataset consists of pairs $(p,a)$, where $p$ represents a programming task description and $a$ represents a corresponding code solution that satisfies all specified test cases. Following conventional practice, we fine-tune the LLM by minimizing the cross-entropy loss, focusing on generating the code solution $a$ given the task description $p$:
@@ 3.3.1. Dataset Balancing and Sampling Strategies
2:$D_{\text{og}}$: Standard instruction tuning dataset
3:$D_{\text{clarify}}$: Clarify-aware instruction tuning dataset
@@
6:$D^{\text{all}}\leftarrow$ Combine $D^{\text{og}}$ and $D^{\text{clarify}}$ according to ratio $r$ and chosen sampling strategy
@@ 5.1. Research Questions
5.2. Dataset (HumanEvalComm V2)
We opted to use HumanEvalComm (Wu and Fard, 2025), a benchmark dataset for evaluating communication skills in LLMs during code generation. HumanEvalComm is built by modifying the widely used HumanEval dataset to include incomplete, inconsistent, and ambiguous requirements. However, after doing a deep dive into the dataset itself, we found a need to make certain corrections - in some cases, the problem description was not modified enough to result in an LLM asking a question, leading to a lower Communication Rate than expected. For those questions, we modified the wording to make sure that they would result in possible clarifying questions. Modifications included rewording certain problems and adding ambiguous, incomplete, or inconsistent variations. The updated dataset, named HumanEvalComm V2, maintains a similar distribution of question types with slight adjustments in some categories. This refined version aims to provide a more robust evaluation of models’ ability to ask clarifying questions. The updated dataset is available in our GitHub repository.
@@ 6.1. Clarification Competency of Code LLMs
Model Comm. Good Q. Pass@1 Test Pass
@@
CodeLlama 4.38% 3.93% 16.08% 34.99%
CodeQwen1.5 Chat 0% 0% 47.56% 62.85%
DeepSeek Chat 28.25% 23.03% 34.76% 42.42%
DeepSeek Coder 24.12% 21.29% 41.95% 56.86%
ClarifyCoder (w/ DS Coder) 57.42% 47.68% 35.3% 47.25%
ClarifyCoder (w/ DS Coder, loss=answer only) 63.61% 51.93% 29.38% 40.96%
Table 3. Results on HumanEvalComm V2 across all clarification categories. First-place results are underlined and bolded, while second-place and third-place results are bolded. The Pass@1 and Test Pass Rate are post-clarification results in the second round to generate code with Q&A information.
Model Pass@1 Test Pass Rate
CodeLlama 29.27% 41.63%
CodeQwen1.5 Chat 76.22% 84.95%
DeepSeek Chat 34.76% 53.00%
DeepSeek Coder 73.62% 82.36%
ClarifyCoder (w/ DS Coder) 73.01% 82.00%
ClarifyCoder (w/ DS Coder,loss=answer only) 74.85% 83.29%
Table 4. Results on HumanEval benchmark for different LLMs in experiments. First-place results are underlined and bolded, while second-place and third-place results are bolded.
@@
To address RQ1, we evaluate the effectiveness of clarify-aware fine-tuning in improving a model’s ability to generate meaningful clarifying questions. We compare our proposed ClarifyCoder with existing code generation models on the HumanEvalComm V2 benchmark, which specifically tests performance on problems requiring clarification.
@@
From Table 3, we observe significant improvements in communication rate with ClarifyCoder (with DeepSeek Coder and answer only). ClarifyCoder achieves a substantially higher communication rate of 63.61%, more than doubling the rate of the original DeepSeek Coder (24.12%) and the second-best model DeepSeek Chat (28.25%). This demonstrates that our technique effectively improves the model’s ability to recognize when clarification is needed. ClarifyCoder similarly excels in generating high-quality clarifying questions, with 51.93% of its questions being evaluated as good, compared to 23.03% for DeepSeek Chat. This indicates that our approach not only increases the frequency of clarifications but also their quality.
Comparing Tables 3 and 4, we observe that all models experience performance degradation when moving from the original HumanEval benchmark to HumanEvalComm, which contains problems requiring clarification. Due to 0% good question rate, CodeQwen1.5 Chat shows the largest relative drop, from 76.22% to 47.56% for Pass@1 (37.6% decrease) and from 84.95% to 62.85% for Test Pass Rate (26.0% decrease). However, the pass rates did not drop further, which indicates a potential data contamination issue for CodeQwen1.5 Chat. DeepSeek Coder experiences a decline from 73.62% to 41.95% for Pass@1 (43.0% decrease) and from 82.36% to 56.86% for Test Pass Rate (30.9% decrease). ClarifyCoder shows a more severe drop for Pass@1 (more than 50% decrease) and for Test Pass Rate (30% decrease).
@@
Interestingly, ClarifyCoder maintains competitive code generation performance while significantly outperforming all other models in clarify-awareness metrics. This suggests that ClarifyCoder effectively enables models to recognize when clarification is needed without sacrificing code generation capabilities. By enhancing the model’s ability to recognize ambiguity and generate relevant clarifying questions, ClarifyCoder demonstrates a more interactive and human-like approach to code generation tasks. These results support our hypothesis that explicit alignment for clarification capabilities can significantly improve model performance on tasks requiring disambiguation, which is crucial for real-world programming scenarios where problem specifications are often ambiguous.
@@
Answer to RQ1: Our experiments demonstrate that clarify-aware fine-tuning substantially improves models’ ability to ask effective clarifying questions, with ClarifyCoder achieving a communication rate of 63.61% and a good question rate of 51.93% (more than doubling baseline models) while maintaining competitive code generation performance.

### v2

# Can Code Language Models Learn Clarification-Seeking Behaviors?

@@ Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities in code generation tasks. However, a significant gap remains between their output and problem-solving strategies of human developers. Unlike human developers, who allocate substantial time to disambiguating requirements through iterative dialogue, LLMs often generate code despite the ambiguities in natural language requriements, leading to unreliable solutions. Different from prior work, we study whether a Code LLM can be fine-tuned to learn clarification-seeking behavior. While recent work has focused on LLM-based agents for iterative code generation, we argue that the fundamental ability to recognize and query ambiguous requirements should be intrinsic to the models themselves, especially in the context of agentic AI where models and humans collaborate frequently. In this research, we present ClarifyCoder, a new framework with synthetic data generation and instruction-tuning that fine-tunes an LLM to identify ambiguities, if they exist, and request clarification before proceeding with code generation. Our approach consists of two main components: (1) a data synthesis technique that augments existing programming datasets with scenarios requiring clarification to generate clarification-aware training data, and (2) a fine-tuning strategy that teaches models to prioritize seeking clarification over immediate code generation when faced with incomplete or ambiguous requirements. We further provide an empirical analysis of integrating ClarifyCoder with standard fine-tuning for a joint optimization of both clarification-awareness and coding ability. Experimental results demonstrate that ClarifyCoder achieves a 63% communication rate (a 40% absolute increase) and a 52% good question rate (30% absolute increase) on ambiguous tasks, significantly improving the communication capabilities of Code LLMs through meaningful clarification dialogues, while maintaining code generation capabilities.
@@ 1. Introduction
Large language models (LLMs) (Vaswani et al., 2017; Svyatkovskiy et al., 2020; Wang et al., 2021; Feng et al., 2020) have demonstrated outstanding capabilities in code generation from natural language requirements, as shown by systems such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022a), and CodeGen (Nijkamp et al., 2022). Given the natural language requirements, or prompt, as input, the LLMs perform code generation that outputs the precise code snippets (Hassan et al., 2024; Mu et al., 2024).
However, unlike programming languages that are designed to be precise and unambiguous, the coding tasks described in natural language are inherently ambiguous (Hoare, 1969). For instance, suppose the coding task is “Return list with elements incremented by a number.” (Wu and Fard, 2025). The description is ambiguous as it does not indicate the specific value of the number. For software engineers, they may ask questions such as “could you clarify the specific value of the number in the requirement?” to disambiguate the description in this coding task. However, several studies  (Wu and Fard, 2025) have shown that LLMs directly generate code rather than asking clarification questions in these cases (Miao et al., 2025; Mu et al., 2024). More specifically, in HumanEvalComm benchmark (Wu and Fard, 2025)–a benchmark for evaluating LLMs’ communication ability, LLMs generate code outputs in over 63% of ambiguous scenarios without seeking necessary clarifications.
The disambiguation of coding tasks is not only necessary but also important to ensure that the generated code is reliable and trustworthy. In real-world problem-solving processes, people, including software engineers, rely on continuous dialog to resolve ambiguities, align mental models, and iteratively refine problem statements and their solutions (Whitehead, 2007; Pressman, 2005; Mistrík et al., 2010). Based on the study in  (Meyer et al., 2019), senior developers allocate only 10–20% to coding, with the majority remainder devoted to code reviews (25%), meetings (20%), and system design (30%). This suggests that developers spend a significant amount of time on task disambiguation, including requirements, design, solutions, and meetings to align the requirements with stakeholders. On the other hand, in the new era of Agentic Software Engineering, the role of intelligent agents is becoming more prominent, where the Software Agents and Human Developers are envisioned to work collaboratively (Hassan et al., 2025). In this context, we anticipate an increasing need and importance for intelligent agents to resolve ambiguities and ensure the trustworthiness of software engineering tasks (Belcak et al., 2025; Hassan et al., 2025) in a human-AI collaboration environment.
To address the above issue, several works have been proposed to disambiguate requirements for code generation using an iterative LLM-based workflow. TICODER (Lahiri et al., 2022) proposes an interactive workflow for guided intent clarification through tests to generate more accurate code. TICODER interacts with the user in the form of formal specifications. ClarifyGPT (Mu et al., 2024) is a multi-step LLM-based code generation workflow that performs intent clarifications using a combination of test generation, code check, reasoning, and refinement. Okanagan (Wu and Fard, 2025) is a multi-round structure with customized prompts for asking clarifying questions when needed in code generation tasks where tasks could be ambiguous. Recently, ClariGen (Miao et al., 2025) has been proposed using three LLMs together with interaction with the user to answer clarifying questions for enhancing code generation.
In spite of continuous research efforts, existing work may still face challenges when applied to code generation, where requirements can be either ambiguous or well-specified. First, user interaction with clarifying questions is still enforced for clear coding tasks that could be completed without clarifications. ClarifyGPT, TICODER, and ClariGen explicitly require the user interaction module, where it is a must that the method needs to interact with the user or LLM-based simulation. Different from the other three methods, Okanagan is set up to either generate code or ask clarifying questions if only clarification is needed (i.e., when the coding task is considered ambiguous). However, Okanagan tends to still ask questions that appear to be unnecessary for original problems that do not need questions (pass@1 dropped from 65% down to 27% for standard coding tasks). This indicates a need to strike a balance between asking unnecessary questions and truly needed questions. Second, all of the existing works use LLM-based workflows and tend to be more complicated and costly than a single LLM. Even if we don’t count the LLM-simulation that provides answer to the clarifying questions, ClariGen and Okanagan needs 3 calls to LLMs with customized prompts for each call, while ClarifyGPT needs 4 calls together with other modules such as test input generation and code consistency checks.
Figure 1. Overview of ClarifyCoder, including the synthetic data construction and clarification-aware fine-tuning.
Our aim in this work is to explore the research question: Can Code LLMs learn clarification-seeking behavior? In the real-world code generation task, the input coding task may be standard or ambiguous. If the input is standard, it may not be necessary to have user interaction each time. We want to understand whether it’s possible to make Code LLMs learn the clarification-awareness behavior that human developers have by changing model weights using techniques such as fine-tuning. To the best of our knowledge, there is little work on empowering Code LLMs to get clarification-aware behaviors for code generation.
@@
In this research, we propose ClarifyCoder, a novel framework that includes synthetic data generation and instruction-tuning techniques to fine-tune a Code LLM to have clarification-seeking behavior. With the goal to enhance a LLM’s clarification awareness in code generation, the framework includes 1) a new synthetic data generation approach for automatically creating ambiguous problem descriptions and corresponding clarifying questions, and 2) a technique combining automated data synthesis with targeted instruction tuning for clarification-aware (or clarify-aware) code generation. As a result, ClarifyCoder can be integrated in the standard fine-tuning that satisfies both clarify-awareness and code generation capabilities.
We evaluate both the utility and clarify-awareness of our clarify-aware fine-tuning based on the proposed synthetic training data generation. Our experiments demonstrate that ClarifyCoder significantly improves LLMs’ ability to recognize ambiguity and generate appropriate clarifying questions, including increases in communication rate from 24% to 63% and good question rate from 21% to 52%, with absolute increases of 40% and 30% respectively, while maintaining the code generation ability. Besides, we also compared ClarifyCoder with training-free methods, including few-shot prompting and CoT reasoning, which showed significant lower clarify-awareness compared to ClarifyCoder. We further break down the results on different clarification categories and find that Incompleteness is easiest for models to detect while Inconsistency is most challenging. We further analyze the perplexity and entropy of the synthetic data generation in ClarifyCoder, and conduct a manual evaluation on LLM-based metrics to investigate the reliance of LLM-based metrics in evaluations. Finally, we investigate the impact of different synthetic data ratios and show that ClarifyCoder could be combined with standard fine-tuning to optimize both clarify-awareness and other utilities.
Our main novelty in the existing literature is that, based on the experiments, the ClarifyCoder fine-tuning technique is able to empower an original LLM to ask questions (60% of the time, more than doubled compared with the original LLM) to an ambiguous (or incomplete, or inconsistent) problem, and generate code (70%+ pass rates) to a normal problem without sacrificing coding performance. Therefore, ClarifyCoder technique, to our knowledge, is one of the first works to achieve this capability in the literature. To advocate open science and full replication, we also released our data, full code on GitHub, and our model on HuggingFace.
Contributions. To summarize, we make the following contributions:
@@
  • We propose a fine-tuning framework, ClarifyCoder, that combines synthetic data generation with targeted fine-tuning to enable clarify-aware code generation for a given LLM. The new synthetic data generation approach automatically generates ambiguous problem descriptions and corresponding clarifying questions.
  • We conduct experiments to evaluate the effectiveness of ClarifyCoder in improving clarify-aware capabilities, while maintaining code generation capabilities of the models. This includes 1) comparison with different LLMs and training-free methods on clarify-awareness, 2) performance breakdown on clarification categories, 3) effectiveness of synthetic data generation, 4) manual evaluation of LLM-based metrics. In addition, we also evaluate the effect of different ClarifyCoder variants such as using different ratios of standard and synthetic data, and calculating loss on answer only vs both problem and answer.
@@
  • A publicly available, open-source repository for full replication and future research on ClarifyCoder, at https://github.com/jie-jw-wu/clarify-coder. The ClarifyCoder model weights are also released at https://huggingface.co/jie-jw-wu/clarify-coder.
@@ 2.1. Backgrounds
  • Maintain the model’s code generation capabilities for well-specified tasks.
  • Enable the model to recognize ambiguity and request clarification when necessary.
@@
Table 1. Comparison of responses of DeepSeek Coder and ClarifyCoder for an ambiguous problem statement. ClarifyCoder is fine-tuned based on DeepSeek Coder and demonstrates the ability to ask clarifying questions to disambiguate coding problems. This example is from an incomplete problem in HumanEvalComm, modified from HumanEval/43 in the HumanEval benchmark.
@@
def candidate(l): """ candidate takes a list of integers as input. it returns True if there are two distinct elements in the list that """ def candidate(l): seen = set() for num in l: if num in seen: return True seen.add(num) return False It seems like your function description is incomplete. Could you please provide more details about what the function should return if there are two distinct elements in the list?
@@ 2.2. Problem Formalization
We adopt the taxonomy of clarification types (ambiguous, inconsistent, and imcomplete problems) from previous works (Wu and Fard, 2025). The taxonomy was designed by following clarification types based on both the literature in Requirement Engineering (RE) and understanding of how feasible can the RE concepts can be applied to problems in HumanEval:
  • Ambiguous Problem: Some statements in the problem descriptions could be ambiguous and correspond to different concepts.
  • Inconsistent Problem: Some statements in the problem descriptions show conflict or inconsistency between each other.
  • Incomplete Problem: Some concepts or conditions are missing in the problem descriptions.
@@
Inspired by (Kobalczyk et al., 2025), we further develop this taxonomy of clarification types into formal definitions for code generation tasks. Let $H^{*}$ denote the ground-truth code solutions that are correct. Let $H$ denote the solution space that comprises all syntactically valid implementations that meet the requirements $R$ of the problem statement. :
@@
Definition 2.1 (Ambiguous Problem (Kobalczyk et al., 2025)).
A coding problem $S$ exhibits ambiguity if $H^{*}\subset H$ ($H^{*}$ is a proper subset of $H$), i.e.,
@@
"""Return list with elements incremented by a number."""
This specification permits multiple valid interpretations $H=\{h_{\text{1}},h_{\text{2}},h_{\text{3}},h_{\text{4}},$...\}. However, the ground truth code solution $H^{*}$ for this problem statement is only a proper subset of $H$. In this example, the ambiguity stems from an underspecified number to increment by. See also another example in Figure 2. Given this ambiguous problem, the goal of this model is to ask clarifying questions to disambiguate the coding problem. When the clarifying question(s) are answered, with this additional information, the model attempts to generate the correct code $h\in H^{*}$.
@@ 3.1. Baseline Instruction Tuning
We first establish a baseline by employing standard instruction tuning on the original programming dataset, denoted as $D^{og}$ . This dataset consists of pairs $(p,a)$, where $p$ represents a programming task description and $a$ represents a corresponding code solution that satisfies all specified test cases. Following conventional practice, we fine-tune the LLM by minimizing the cross-entropy loss, focusing on generating the code solution $a$ given the task description $p$:
@@ 3.3.1. Dataset Balancing and Sampling Strategies
[…]
@@
[…]
@@ 5.1. Research Questions
5.2. Dataset (HumanEval and HumanEvalComm)
We tested different models on coding tasks with both clear requirements and ambiguous requirements. For clear requirements, we used the widely used HumanEval benchmark. For ambiguous requirements, we opted to use HumanEvalComm (Wu and Fard, 2025), a benchmark dataset for evaluating communication skills in LLMs during code generation. HumanEvalComm is built by modifying the widely used HumanEval dataset to include incomplete, inconsistent, and ambiguous requirements. One of the authors in this paper, did a thorough exam on the dataset to make corrections if needed. We found that in a small number of cases, the problem description was not modified enough to result in an LLM asking a question, leading to a lower Communication Rate than expected. For those questions, we modified the wording to make sure that they would result in possible clarifying questions. Modifications included rewording certain problems and adding ambiguous, incomplete, or inconsistent variations. The updated dataset maintains a similar distribution of question types with slight adjustments in some categories. This refined version aims to provide a more robust evaluation of models’ ability to ask clarifying questions. The updated dataset is checked into the HumanEvalComm dataset on GitHub (Wu and Fard, 2025) and HuggingFace (hum, 2024) after careful verifications.
@@ 6.1. Clarification Competency of Code LLMs
Model HumanEvalComm HumanEval
Comm. Good Q. Pass@1 Test Pass
@@
CodeLlama 4.38% 3.93% 29.27% 41.63%
CodeQwen1.5 Chat 0% 0% 76.22% 84.95%
DeepSeek Chat 28.25% 23.03% 34.76% 53.00%
DeepSeek Coder 24.12% 21.29% 73.62% 82.36%
ClarifyCoder (w/ DS Coder) 57.42% 47.68% 73.01% 82.00%
ClarifyCoder (w/ DS Coder, loss=answer only) 63.61% 51.93% 74.85% 83.29%
Table 3. Results on HumanEvalComm and HumanEval benchmarks for different LLMs. First-place and Second-place results are underlined and bolded, while top 4 results are bolded. ”loss=answer only” denotes the loss during fine-tuning, calculated only on the answers rather than on both the problems and the answers.
(a) Multi-dimensional performance profile for clear (HumanEval) and ambiguous requirements (HumanEvalComm).
(b) a scatter plot of the trade-off between clarification-seeking ability and code generation performance. Dashed arrow represents the impact of Clarification-aware fine-tuning on the original DeepSeek Coder model.
Figure 3. Comprehensive performance analysis of ClarifyCoder and baseline LLMs. ”loss=answer only” denotes the loss during fine-tuning, calculated only on the answers rather than on both the problems and the answers.
Table 4. Results of different methods on HumanEval benchmark (well-specified coding tasks). Pass@1 and Test PR measure code generated in the first (only) round. The top two results are bolded. Pass rates are considered unavailable for methods where user interaction to perform clarification is enforced.
Method Method Type # LLM calls Interaction Enforced Pass@1 Test Pass Rate
TICODER LLM Workflow 2 Yes N/A N/A
ClarifyGPT LLM Workflow 4 Yes N/A N/A
ClariGen LLM Workflow 3 Yes N/A N/A
Okanagan(Base=ChatGPT) LLM Workflow 3 No 27.45% 33.45%
Okanagan(Base=DS Coder) LLM Workflow 3 No 21.25% 24.3%
DeepSeek Chat Single LLM 1 No 34.76% 53.00%
DeepSeek Coder Single LLM 1 No 73.62% 82.36%
ClarifyCoder (w/ DS Coder) Single LLM 1 No 73.01% 82.00%
ClarifyCoder (w/ DS Coder, loss=answer only) Single LLM 1 No 74.85% 83.29%
@@
To address RQ1, we evaluate the effectiveness of clarify-aware fine-tuning in improving a model’s ability to generate meaningful clarifying questions. We compare our proposed ClarifyCoder with existing code generation models on the HumanEvalComm benchmark, which specifically tests performance on problems requiring clarification.
@@
Table 3 displays the numerical results on HumanEvalComm and HumanEval benchmarks for different LLMs and ClarifyCoder. We included two variants of ClarifyCoder: one where loss during fine-tuning is calculated on both the problem and answer, and another one, with donation ”loss=answer only”, where loss is calculated solely on the answer. From Table 3, we observe significant improvements in communication rate with both variants of ClarifyCoder. ClarifyCoder achieves a substantially higher communication rate of 63.61% (for with answer only), more than doubling the rate of the original DeepSeek Coder (24.12%) and the second-best model DeepSeek Chat (28.25%). This demonstrates that our technique effectively improves the model’s ability to recognize when clarification is needed. ClarifyCoder similarly excels in generating high-quality clarifying questions, with 51.93% of its questions being evaluated as good, compared to 23.03% for DeepSeek Chat. This indicates that our approach not only increases the frequency of clarifications but also their question quality.
@@
Figure 3(a) shows a radar chart that shows each model’s capability on four metrics regarding coding performance and clarification-seeking ability. Figure 3(b) shows a scatter plot of the fundamental trade-off relationship between asking quality clarifying questions and generating correct code, with most models clustered in suboptimal regions. Comparing the multi-dimensional performance profile and the performance trade-off visualization in Figure 3, we observe that the two variants of ClarifyCoder experience very small performance degradation for clear requirements on HumanEval benchmark, while showing significant improvements in communication rates and Good Question Rates for ambiguous requirements on HumanEvalComm benchmark. This represents a good trade-off and advantage for adopting our clarification-aware fine-tuning. This suggests that ClarifyCoder effectively enables the models (in our case, DeepSeek Coder) to recognize when clarification is needed without sacrificing code generation capabilities. By enhancing the model’s ability to recognize ambiguity and generate relevant clarifying questions, ClarifyCoder demonstrates a more interactive and human-like approach to code generation tasks. These results support our hypothesis that explicit alignment for clarification capabilities can significantly improve model performance on tasks requiring disambiguation, which is crucial for real-world programming scenarios where problem specifications are often ambiguous.
It should be noted that in this work, we focused on Comm. Rate and Good Question Rate in HumanEvalComm. Comm. Rate and Good Question Rate (in the first round) are indeed much more important metrics than the test pass rates and Pass@1 (in the second round). The reason is that: if an ideal LLM is given an ambiguous (or incomplete, or inconsistent) coding problem, and is instructed to ask questions or generate code, the LLM must ask questions ideally. The above results show that, to our knowledge, ClarifyCoder is one of the first models in the literature that is able to intelligently ask questions and know when not to ask questions. Table 4 summarizes differences between our models and existing methods, such as Okanagan (Wu and Fard, 2025), TICODER (Lahiri et al., 2022), ClarifyGPT (Mu et al., 2024), ClariGen (Miao et al., 2025), which enforce user interaction for clarification. In contrast, ClarifyCoder exhibits a unique clarification-seeking behavior that more closely resembles human interaction.
@@
Answer to RQ1: Our experiments demonstrate that clarify-aware fine-tuning substantially improves the original model’ ability to ask effective clarifying questions, with ClarifyCoder achieving a communication rate of 63.61% and a good question rate of 51.93% (more than doubling the original model, i.e., DeepSeek Coder) while maintaining competitive code generation performance.